OpenStudy (anonymous):

Which is a counterexample to this conjecture? The sum of any two consecutive integers is a composite number. A. 16 + 17 = 33 B. 10 + 11 = 21 C. 6 + 7 = 13 D. 7 + 8 = 15

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

I am realy realy bad at math.

OpenStudy (turingtest):

try to find two consecutive integers that add to a prime number

OpenStudy (anonymous):

what is composite number?

OpenStudy (anonymous):

A number that has one and itself.

OpenStudy (anonymous):

what about 13?

OpenStudy (anonymous):

Does 13 have one and itself as multiples.

OpenStudy (anonymous):

33=1*3*11 21=1*3*7 13=1*13 15=1*3*5 1 is factor of every number. so you can know.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \rm Composite~~numbers }\), are the opposite of prime numbers. \(\normalsize\color{blue}{ \rm Composite~~numbers }\): \(\normalsize\color{blue}{ \rm 4 }\) (can be divisible by 2) \(\normalsize\color{blue}{ \rm 6 }\) (can be divisible by 2 & 3) \(\normalsize\color{blue}{ \rm 8 }\) (can be divisible by 2 & 4) \(\normalsize\color{blue}{ \rm 9 }\) (can be divisible by 3) \(\normalsize\color{blue}{ \rm 10 }\) (can be divisible by 2 & 5) and etc., \(\normalsize\color{red}{ \rm Prime~~numbers }\) would be starting like: \(\normalsize\color{red}{ \rm 2 }\), \(\normalsize\color{red}{ \rm 3 }\), \(\normalsize\color{red}{ \rm 7 }\), \(\normalsize\color{red}{ \rm 11 }\) and \(\normalsize\color{red}{ \rm on... }\)

OpenStudy (anonymous):

yes 13 is a composite number.

OpenStudy (anonymous):

I mean prime sorry.

OpenStudy (solomonzelman):

yes, it prime

OpenStudy (anonymous):

prime numbers have one and itself as multiples.

OpenStudy (solomonzelman):

\(\normalsize\color{red}{ \rm prime~~number }\) can be only divided by \(\normalsize\color{red}{ \rm itself }\) and by \(\normalsize\color{red}{ \rm 1 }\). Like, \(\normalsize\color{red}{ \rm 13 }\) can be only divided by \(\normalsize\color{red}{ \rm 13 }\) (itself), or by \(\normalsize\color{red}{ \rm 1 }\) .

OpenStudy (solomonzelman):

do you understand the conjecture?

OpenStudy (solomonzelman):

or not?

OpenStudy (anonymous):

I understand.

OpenStudy (anonymous):



OpenStudy (solomonzelman):

just to make sure, which answer choice did you choose?